The sum of the measures of two supplementary angles is 180 degrees. If one angle measures 45 degrees less than twice the measure of its supplement, find the measure of each angle. (smallest first)
Answer by algebrahouse.com(1659)

x = one angle
2x - 45 = other angle {one angle is 45 less than twice its supplement}

x + 2x - 45 = 180 {supplementary angles add up to 180°}
3x - 45 = 180 {combined like terms}
3x = 225 {added 45 to each side}
x = 75 {divided each side by 3}
2x - 45 = 105 {substituted 75, in for x,into 2x - 45}

75° and 105° are the two supplementary angles

For more help from me, visit: www.algebrahouse.com